### v1

# TransitReID: Transit OD Data Collection with Occlusion-Resistant Dynamic Passenger Re-Identification

Abstract
Transit Origin-Destination (OD) data are essential for transit planning, particularly in route optimization and demand-responsive paratransit systems. Traditional methods, such as manual surveys, are costly and inefficient, while Bluetooth and WiFi-based approaches require passengers to carry specific devices, limiting data coverage. On the other hand, most transit vehicles are equipped with onboard cameras for surveillance, offering an opportunity to repurpose them for edge-based OD data collection through visual person re-identification (ReID). However, such approaches face significant challenges, including severe occlusion and viewpoint variations in transit environments, which greatly reduce matching accuracy and hinder their adoption. Moreover, designing effective algorithms that can operate efficiently on edge devices remains an open challenge. To address these challenges, we propose TransitReID, a novel framework for individual-level transit OD data collection. TransitReID consists of two key components: (1) An occlusion-robust ReID algorithm featuring a variational autoencoder guided region-attention mechanism that adaptively focuses on visible body regions through reconstruction loss-optimized weight allocation; and (2) a Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient and robust transit OD matching which balances storage, speed, and accuracy. Additionally, a multi-threaded design supports near real-time operation on edge devices, which also ensuring privacy protection. We also introduce a ReID dataset tailored for complex bus environments to address the lack of relevant training data. Experimental results demonstrate that TransitReID achieves state-of-the-art performance in ReID tasks, with an accuracy of approximately 90% in bus route simulations.
Index Terms:
Person Re-identification, Origin-Destination Collection, Computer Vision, Intelligent Transportation
I Introduction
Transit Origin-Destination (OD) data collection plays a critical role in the fields of transportation engineering for urban planning, service frequency adjustments, and route optimization. It aids in the analysis of relationships and impacts among regional economic development, infrastructure, and urban mobility [1]. This data represents the demand of passengers and is needed to design the bus routes efficiently to minimize travel time, reduce operating costs, and driver scheduling. As shown in Fig. 1, traditional methods for estimating transit OD data rely on manual approaches such as surveys, which are labor-intensive and suffer from low response rates [2].
More advanced methods, such as those utilizing mobile phone data[3] and Bluetooth technology [4], require passengers to carry specific devices with WiFi or Bluetooth enabled, which limits coverage and might raise privacy concerns due to the unique identifiers of the device. Automated Passenger Counters (APCs) [5] also assist in estimating OD data, however, they can only capture the aggregate number and locations of boarding and alighting passengers. They fail to match passengers at the individual level, which is crucial for studies focused on individual passenger behavior.
Meanwhile, most transit vehicles in the US are already equipped with onboard cameras for liability and surveillance purposes, presenting a valuable opportunity to repurpose them as smart sensors for transit OD data collection. By integrating these cameras with person re-identification (ReID) techniques and edge computing devices, seamless individual-level matching can be achieved [6]. By integrating sensors such as GPS, the matching results can be readily translated into transit OD data.
Despite the advantages of camera-based methods over prevailing techniques, their adoption in transit OD data collection remains limited due to challenges posed by the transit vehicle environment, particularly occlusion issues. For example, during peak times, boarding and alighting passengers may obscure each other from the camera view, leading to information loss and background confusion. Recent research increasingly focuses on occlusion issues in ReID, broadly categorized into two types: methods based on local attention mechanisms and those based on feature recovery. The former constructs attention mechanisms by focusing on different local areas to suppress interference from occluded parts [7], while the latter attempts to recover information from occluded areas by associating temporal or spatial information [8, 9]. Fixed local regions are used to differentiate various body parts of pedestrians and assign scores to different areas using a quality predictor in [10]. However, this fixed-area approach is not suitable for passenger matching tasks on transit vehicles under dynamic viewpoints. As the distance and angles between passengers and cameras change over time, the regions of their body parts shift throughout the video.
Figure 1: Traditional methods for estimating transit OD data and our method.
Due to limitations in power consumption, memory, and computational resources, edge devices are generally incapable of executing advanced and complex analytical algorithms[11, 12]. Fast and accurate estimation of OD (origin-destination) data is essential for enabling real-time passenger flow analysis in transit systems. However, most deep learning algorithms have relatively long inference times on edge devices, leading to latency and performance degradation. Some efforts have focused on model lightweighting techniques, such as knowledge distillation and pruning, or on optimizing neural network architectures specifically for edge hardware[13]. Nevertheless, lightweight models often suffer from a trade-off in performance, with reduced accuracy to some extent.
What’s more, unlike traditional ReID studies, our work focuses on the problem of transit OD data collection, hence the gallery, shown in Fig. 2, is dynamic, changing with the variations in passengers boarding and alighting at each station. Consequently, ReID evolves from a problem of matching 1 query to $N$ options, into a dynamic 1 to $N_{1},N_{2},...N_{K}$ problem, where $N_{k}$ represents the size of the gallery updated at station $k$. To address this, we design a hierarchical storage and dynamic matching (HSDM) method to balance the trade-offs among matching accuracy, storage pressure, and processing speed.
To jointly tackle the aforementioned issues of occlusion and the constraints of online adaptive matching, we propose TransitReID, an individual-level and robust transit OD data collection method that incorporates occlusion-resistant and dynamic passenger ReID. The overall framework is shown in Fig. 3, featuring two principal technical innovations: (1) An occlusion-robust ReID algorithm that employs a region-attention module to adaptively weight less obscured body segments with clearer viewpoints, guided by a co-trained Variational Autoencoder (VAE) grading model. The VAE’s reconstruction loss directly informs attention weight allocation, emphasizing visually coherent regions. This is followed by a feature fusion strategy that preserves discriminative information through concatenated weighted averages of region embeddings. (2) A Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient transit OD matching, which incorporates Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. A multi-threaded design is implemented to support near real-time operation on transit vehicles equipped with edge devices, without requiring any modifications to the models or architectures. Notably, the framework processes video locally on edge devices and stores only high-level, non-identifiable features, ensuring a privacy-preserving and non-intrusive solution.
Another issue is that current video ReID datasets [14, 15, 16, 17] are not specifically designed for occlusion scenarios, while most occluded person ReID datasets [18, 19, 20, 21] are image-based and not suited for training tasks under transit surveillance video conditions. Additionally, person images in those datasets are captured from consistent angles, whereas in transit scenarios, the angles of passenger images captured by surveillance cameras change continuously. This variability is crucial for training a robust and generalizable model under transit bus operational scenarios. Therefore, we construct a passenger ReID dataset collected from real bus operational settings, including footage from cameras at the front and rear doors. This dataset captures 17,636 images of 157 passengers.
In essence, the major contributions of this work are as follows:
  • Onboard transit surveillance systems are combined with computer vision technology to collect transit OD data, achieving individual-level precision in the construction of the OD matrix.
  • An attention-based passenger ReID algorithm is specifically designed for transit vehicles scenarios with variable occlusion and viewpoints.
  • A hierarchical storage and dynamic matching (HSDM) mechanism tailored for actual transit scenarios is designed, transforming the traditional static ReID problem into a dynamic ReID challenge.
  • An efficient programming approach is designed to enable real-time operation on edge devices for smart transit applications. In addition, edge-based processing inherently supports privacy protection.
Figure 2: Segmentation results after detection and tracking.
II Related Work
II-A Transit OD Data Collection
Transit OD data collection refers to the systematic process of gathering information about the starting points (origins) and endpoints (destinations) of individual trips within a public transportation network. The most traditional and ubiquitous methods for estimating transit OD data is by surveys [22, 23], which are labor-intensive and have limited response rates. In [24], an integrated program based on Automated Fare Collection (AFC) systems was developed that can generate OD matrices and passenger load profiles by making some assumptions about round trips. A trajectory search algorithm was proposed to process passenger smart card data to track daily travel trajectories in [25]. An optimization model was proposed in [26] to approximate network-level OD flows with big data generated from Automated Passenger Counters (APCs). RFID technology was utilized to create a Boarding-In and Boarding-Out (BIBO) system for passenger monitoring in [27].
Moreover, mobile phone data [3], Bluetooth technology [4], and multi-source approaches – such as combining GIS data with passenger IC card data to construct OD matrices, as demonstrated in [28] – have been utilized. However, these methods face challenges, including high costs, significant maintenance requirements [2], and the need for passengers to carry specific devices, which limits data coverage. Additionally, the extensive use of these systems in public transport can lead to privacy concerns [29], such as exposing details about ticket types, passenger categories, travel behaviors, and purchasing methods.
While most methods capture only the number and locations of boarding and alighting passengers, they lack the precision to individually track and match passengers, crucial for detailed behavioral studies. Video-based passenger re-identification (ReID) technology combined with transit surveillance cameras provides a potential solution to this problem. In [6], the YOLO algorithm is utilized to detect passengers, generating feature vectors through a neural network and selecting matches based on vector distance, but this work ignores the passive influence of occlusion. Several works [30, 31, 32] address the problem by only collecting information from head images of passengers inside the bus. However, this approach overlooks the potential rich information from other regions, lacks discrimination power against similar headgear, and is not robust to changes in lighting.
Our proposed methodology not only maximizes the useful information extracted from each passenger through an attention mechanism but also enhances privacy protection by processing data locally on onboard edge devices, ensuring that only high-level features are stored.
II-B Occluded Re-Identification
Occlusion is by far the major problem in the computer vision-based person ReID task. Most of the occluded person re-identification methods rely on convolutional neural networks (CNNs)[7]. These aim at body part-level feature measurement, for example, a region-based quality predictor was proposed and trained with the regional feature generation model to get the weighted feature representation in [10]. However, the regions defined in this work is fixed, thus not suitable for dynamic perspectives. A pose-guided feature alignment (PGFA)[18] utilized human pose landmarks to build an attention map, guiding the model to extract useful information from non-occluded parts. Similarly, [33, 34, 35, 36] utilize posture-guided attention mechanisms to discern visible areas. These methods partition the ReID problem into two segments: initially detecting key areas, then selectively extracting regional features for matching. However, the accuracy of the posture estimation model can directly impact subsequent discrimination. Other methods include feature recovery[8, 9, 37, 38, 39], graph convolution[40], and multi-model methods[41, 42]. Unfortunately, these methods often struggle with generalization across diverse scenarios, such as changes in viewpoints, variations in environmental lighting, and complex occlusions among passengers in transit vehicle environments. Additionally, the lack of scenario-specific training data further limits the models’ generalizability. Moreover, they address static ReID problems, whereas the gallery in a transit vehicle operating environment is inherently dynamic. We design a hierarchical storage and dynamic matching (HSDM) mechanism specifically adapted for dynamic ReID, detailed in Section III. F.
Figure 3: The Framework of the Proposed TransitReID. In Dynamic Transit ReID, the surveillance system captures image sequences of the same passenger boarding and alighting at different stops. These images are processed by a feature generation network to extract one-dimensional feature representations. If the passenger is boarding, the features are temporarily stored in the gallery. If the passenger is alighting, the system retrieves a matching result from the gallery through the hierarchical storage and dynamic matching (HSDM) mechanism. In the Feature Generation Network, the input consists of an image sequence of a passenger. After action detection and segmentation, each mask is processed by a human parsing network, which divides the image into three body parts: ‘head’, ‘torso’, and ‘legs’. Each part is then fed into a grading model and a baseline model to obtain a quality score and an initial feature representation. Finally, a feature averaging unit integrates these outputs into the final 3*2048 length feature vector.
III Methodology
III-A Problem Definition
Person ReID is a computer vision problem for matching the same person across images captured by multiple heterogeneous cameras. In public transportation scenarios, surveillance systems in transit vehicles play a crucial role in enhancing vehicle safety and supporting operational management. However, current systems have not yet expanded to include additional artificial intelligence applications to maximize their performance. The challenges of passenger ReID in transit vehicle scenarios include temporal and spatial variations among different cameras, such as changes in lighting and varying shooting angles. The matching process is further disrupted by occlusions between passengers and between passengers and objects.
In the following sections, we focus on the transit bus environment, which is equipped with two cameras: a front-door camera $C_{f}$ and a rear-door camera $C_{r}$, which capture sequences of images over a time interval $T$. The image sets captured by each camera are:
$I_{f}=\{i_{f,1},i_{f,2},...,i_{f,T}\},\quad I_{r}=\{i_{r,1},i_{r,2},...,i_{r,T}\}.$
Each image may contain multiple passengers. The gallery, $G$ consists of all passenger images extracted from the raw images recorded by both cameras while onboard the bus:
$G=G_{f}\cup G_{r}=\{g_{f,1},g_{f,2},...,g_{f,M_{f}}\}\cup\{g_{r,1},g_{r,2},...,g_{r,M_{r}}\},$
where $M_{f}$ and $M_{r}$ are the number of on board passengers captured by the front and rear cameras, respectively, and $M=M_{f}+M_{r}$ is the total gallery size.
At each bus stop, the query set, $Q$, consists of passengers departing from either the front or rear door:
$Q=Q_{f}\cup Q_{r}=\{q_{f,1},q_{f,2},...,q_{f,N_{f}}\}\cup\{q_{r,1},q_{r,2},...,q_{r,N_{r}}\},$
where $N_{f}$ and $N_{r}$ are the number of passengers alighting from the front and rear doors, respectively, and $N=N_{f}+N_{r}$ is the total number of alighting passengers.
Each passenger $p_{k}$ is mapped to a high-dimensional feature space via a function:
$\phi:p_{k}\rightarrow\mathbb{R}^{d},$
where $d$ is the feature space dimension. For each alighting passenger $q_{k}\in Q$, we determine the best match from the gallery $G$ by computing the distance between $\phi(q_{k})$ and $\phi(g_{j})$ for all $g_{j}\in G$, selecting the closest match:
$\mathbf{min}\ distance(\phi(q_{k}),\phi(g_{j})),\quad\forall g_{j}\in G.$
III-B Detection and Tracking
The video footage recorded by the surveillance cameras at the front and rear doors of the bus is used for the research in this work. Firstly, passengers boarding and alighting are detected with YOLO11 and integrated with NVIDIA TensorRT to speed up the processing. Similar to our previous work [43], $R_{door}$ and $R_{inside}$ are defined in the images to respectively represent the areas outside and inside the bus door based on prior information, and the behavior of passengers is determined by the intersection and union logic of the passenger bounding box centers with the ROIs. Fig. 4 shows a case of a passenger alighting. The behavior classifier is defined with a rule-based algorithm:
Define the starting point as $S=(x_{s},y_{s})$ and ending point as $E=(x_{e},y_{e})$ for a trajectory, there are four possible conditions:
$\big{(}S\in R_{\text{inside}}\big{)}\land\big{(}S\notin R_{\text{door}}\big{)}\land\big{(}E\in R_{\text{door}}\big{)}$ (1)
$\big{(}S\in R_{\text{door}}\big{)}\land\big{(}E\in R_{\text{inside}}\big{)}\land\big{(}E\notin R_{\text{door}}\big{)}$
$\big{(}S\in R_{\text{inside}}\big{)}\land\big{(}S\notin R_{\text{door}}\big{)}\land\big{(}E\in R_{\text{inside}}\big{)}\land\big{(}E\notin R_{\text{door}}\big{)}$
$\big{(}S\in R_{\text{door}}\big{)}\land\big{(}E\in R_{\text{door}}\big{)}$
These equations correspond to trajectories of alighting, boarding, moving while inside the bus, and moving while remaining outside. Only trajectories satisfying the first two cases are retained, with their corresponding locations recorded as key-value pairs.
It has been proven in our previous work that using segmented images can improve the accuracy of ReID, so the same processing strategy is adopted here. Fig. 2 shows some passenger images after body segmentation.
Figure 4: One alighting example. The green box represents the outdoor area and the red box represents the area inside the bus. The blue trajectory shows the moving of the passenger.
III-C Data Preprocessing
Occlusion within transit vehicles significantly impacts the accuracy of conventional ReID methods, especially during peak times when multiple passengers are boarding or alighting. This can lead to information loss and background interference. The target passenger may be disturbed by non-passenger occlusions, e.g., seats or luggage, and non-target passengers [44]. Using the bounding box of the current target passenger as input can result in the loss of passenger information, such as an occluded torso, and may also introduce interference, such as mistakenly including parts of other passengers’ bodies or objects, leading to feature confusion.
The segmentation algorithm mentioned in the previous section addresses the problem of background interference. To address the matching confusion caused by missing information, the global feature representation is abandoned in favor of local feature representations of different body parts. The method from [45] is employed, utilizing the Pascal-Person-Part dataset [46] which segments each passenger’s body image into 7 parts. In this work, these seven parts are aggregated into three main parts according to conventional logic: Head, Torso, and Legs. ‘Torso’ comprises ‘Torso’, ‘Upper Arms’, and ‘Lower Arms’, while ‘Legs’ includes ‘Upper Legs’ and ‘Lower Legs’. Each partial image is converted to gray scale image and cropped according to the minimum bounding box, then resized to $128\times 128$.
III-D Grading Model
Passengers’ local features should be weighted differently to account for various occlusion and viewpoint scenarios. Existing work [10] utilizes a convolutional network including a convolution layer and a fully connected layer to compute local scores. [47] designs a spatial-patch contrastive (SPC) loss to compute aggregation maps to focus on different areas. Our grading model uses a Variational Autoencoder (VAE) to assign scores to images of different body parts, where a higher score indicates better image quality, with less occlusion and a clearer viewpoint. This score will be used to calculate image weights in the subsequent processing stages.
Figure 5: Structure of the Variational Autoencoder (VAE) in the grading model.
The structure of the grading model is shown in Fig. 5. The image input passes through an encoder, which consists of a convolutional layer, followed by a max-pooling layer after each convolution. The features are then processed by residual modules to obtain a flattened feature map. Two fully connected layers calculate the mean and variance of the latent space, introducing randomness to the model and enabling better learning of reconstruction features. The decoder consists of a series of transposed convolutional layers and residual modules which output the reconstructed image through a sigmoid function.
During the training, the grading model is jointly trained with the feature extraction network Global-guided Reciprocal Learning (GRL) [48]. The underlying logic is straightforward: through training, the VAE is enabled to better reconstruct images of higher quality and poorly reconstruct those with severe occlusions and bad viewpoints. The image score, $I_{S}$, is calculated based on the joint similarity between the reconstructed and original images:
$I_{S}=\sigma_{1}SSIM+\sigma_{2}HS+\sigma_{3}MSE$ (2)
To comprehensively assess the image reconstruction effect, the above formula considers three measurement factors. The Structural Similarity Index Measure (SSIM) [49] is an indicator to measure the visual similarity of two images based on three comparison parameters: luminance, contrast, and structure. Histogram Similarity (HS) focuses on the pixel value distribution of images, calculating the similarity of the histograms of the raw and reconstructed images. The Mean Squared Error (MSE) reflects the pixel value differences between the two images. The score is then fed into the Score Mapping for further process.
Score Mapping
A Logarithmic Transformation is used to map the original scores to a range from 0 to 1. Additionally, the logarithmic transformation compresses a wide score range into a smaller output range, which benefits our scenarios, where severe occlusions are of greater concern, the mapped scores decrease rapidly, helping to effectively suppress high-occlusion situations. For cases with minor occlusions, we aim to preserve the features of these images to the greatest extent. Equation (3) defines a mapping $\mathcal{L}:x\rightarrow y$, where $x$ is the input score and $y$ is the transformed score. $min/max$ is the upper/lower bound of the input. Fig. 6 displays the mapping curves at different values of $k$.
$y=\frac{\log(1+k\cdot(x-\min))}{\log(1+k\cdot(\max-\min))}$ (3)
Figure 6: Logarithmic transformation curves with different $k$ values.
III-E Selective Quality Feature Averaging
Initially, a set of images of a passenger is extracted from a video, $\{I_{1},I_{2},...,I_{l}\}$, where $l$ denotes the length of the image sequence. These images are sequentially input into a feature extraction network, which outputs a fixed-length one-dimensional vector as the feature description of the passenger. We employ GRL[48] as the feature extraction network. It effectively utilizes both global and local feature correlations within video frames. This Global-guided Reciprocal Learning (GRL) framework incorporates two key modules: the Global-guided Correlation Estimation (GCE) and the Temporal Reciprocal Learning (TRL). The GCE module generates correlation maps to distinguish between high- and low-correlation features under global guidance, enhancing the model’s focus on relevant features. The TRL module optimizes the use of these features over time, improving the discrimination of temporal information.
The features for three body parts Head, Torso, Legs ($H,T,L$), are processed separately to obtain the feature sets $\mathfrak{F}=\{F_{H},F_{T},F_{L}\}$, where $\underset{k=H,T,L}{F_{k}}=\{f_{1}^{k},f_{2}^{k},...,f_{l}^{k}\}$. The score sets output by the grading model are denoted by $\mathcal{S}=\{S_{H},S_{T},S_{L}\}$, where $\underset{k=H,T,L}{S_{k}}=\{s_{1}^{k},s_{2}^{k},...,s_{l}^{k}\}$. We use $l=8$ frames per person for consistency.
Figure 7: Selective Quality Feature Averaging. Each image is attached with a score reflecting its degree of occlusion. Averaging all frames would suppress those of high quality. Averaging only frames with scores above a threshold, however, would preserve the features of high quality.
Unlike the baseline model [48], we do not consider all the images in a sequence. Continuing to perform weighted averaging across all images could suppress features from images containing significant information, potentially leading to confusion when identifying individuals who have similar local body features but are distinct persons. To avoid this, only images with scores exceeding a certain threshold are retained, thereby preserving the contribution of high-quality images to the final feature, as depicted in Fig. 7. Ultimately, the feature vector for a passenger is calculated by the following formula:
$\mathcal{F}=\underset{k=H,T,L}{\mathtt{CONCAT}}\;(\frac{F_{k}\odot S_{k}}{l-d})$ (4)
Where $d$ is the number of images in the sequence whose scores fall below a certain threshold. $\odot$ is the Hadamard product. $\mathtt{CONCAT}$ indicates that vectors from the three parts are concatenated along the last dimension to form a one-dimensional feature vector. These features would then be stored in the Facebook AI Similarity Search (FAISS) library [50] which is an optimized library for large-scale high dimensional nearest neighbor similarity search.
III-F Hierarchical Storage and Dynamic Matching
Unlike traditional ReID problems, the operational scenario in transit vehicles is dynamic and spatiotemporally continuous. This means that the passenger list onboard updates dynamically over time, resulting in a continuously changing gallery. Following normal logic, passengers in the gallery who have been matched should be removed, which not only reduces the number of samples, decreasing the error rate of subsequent matches, it also facilitates memory savings, which is crucial for edge devices with limited computing power and storage. However, since the matching accuracy is not 100%, directly deleting each match is not advisable as it could lead to error accumulation. Therefore, we propose a hierarchical storage and dynamic matching (HSDM) method to optimize the ReID problem under actual operating scenarios, with its structure shown in Fig. 8.
Figure 8: Structure of hierarchical storage and dynamic matching.
The HSDM consists of 3 major steps as follows, and the processing logic is summarized in Algorithm 1.
III-F1 Hot/Cold Matching and Storage
The passengers are processed sequentially by bus stop in order $1\rightarrow 2\rightarrow...\rightarrow S$. The features of passengers boarding at each stop are added to the FAISS index to form a gallery. For each alighting passenger, we retrieve the Top-$K$ candidates (in our setting $K=5$) from FAISS, obtaining the gallery IDs from rank-1 to rank-K along with their corresponding distances. Then, the match confidence $\gamma$ is calculated based on the distance ratio between rank-1 and rank-2:
$\gamma=1-dist_{rank1}/dist_{rank2}$ (5)
A confidence threshold $\sigma$ is selected based on the distribution of feature distances of positive (correct) and negative (incorrect) samples. If $\gamma>\sigma$, the match is considered successful (Hot Storage), and the matched gallery ID is deleted from FAISS. If $\gamma\leq\sigma$, it is considered a temporary match (Cold Storage), and the gallery ID is not deleted. Instead, it is recorded that the alighting passenger’s temporary match is the rank-1 gallery ID, along with the remaining alternative gallery IDs (rank-2 to rank-5).
III-F2 Snatch
If at the next stop or in subsequent processing, another new alighting passenger also matches a gallery ID that is still in temporary matching, and the confidence is higher, this match will snatch that gallery ID. The ID will be removed from the original alighting passenger’s temporary match and matched to the snatcher, and depending on the level of confidence, it will be confirmed as a successful match (and thus removed from FAISS) or remain a temporary match. For the passenger who is snatched from, they will need to take another ID from previously stored rank-5 alternatives to attempt a match.
III-F3 Final Update
After all stations have been processed, passengers who have not been matched and are still in a temporary matching state automatically transition to a successful match, and the matching accuracy for each station is then calculated.
Algorithm 1 HSDM: Overall Process
1:(Hot/Cold Matching and Storage)
2:for $i=1$ to $S$ do
3:     for each passenger $p_{\text{board}}$ boarding at station $i$ do
4:         Add $p_{\text{board}}.\text{feature}$ to FAISS index
5:     for each passenger $p_{\text{alight}}$ alighting at station $i$ do
[…]
7:         Calculate $\gamma$
8:         if $\gamma>\sigma$ then
9:              Hot Storage:
[…]
12:         else
13:              Cold Storage:
[…]
15:              $p_{\text{alight}}.\text{alt\_IDs}=\text{topK\_IDs}[:K]$
16:(Snatch)
17:for each new passenger $p_{\text{alight2}}$ at station $j>i$ (if exists) do
[…]
19:     Calculate $\gamma_{2}$
[…]
21:         if $\gamma_{2}>\gamma_{\text{old}}$ then
[…]
23:              if $\gamma_{2}>\sigma$ then
[…]
26:              else
[…]
28:                  $p_{\text{alight2}}.\text{alt\_IDs}=\text{topK\_IDs2}[:K]$
29:              $p_{\text{alight}}$ tries remaining alt_IDs and recomputes confidence
30:         else
31:              $p_{\text{alight2}}$ tries remaining alt_IDs and recomputes confidence
32:(Final Update)
33:for each $p_{\text{alight}}$ still in Cold Storage and not snatched do
[…]
35:Compute final matching accuracy
Figure 9: OD data collection in 10 bus stops. The first and last rows represent the 10 boarding and alighting stops, while the middle row corresponds to 60 passengers. Gray lines indicate correct matches, whereas green lines represent incorrect matches.
Figure 10: Top 5 retrieved results in the simulation with 10 bus stops. The target passengers getting off/on the bus are marked with red/green boxes. From left to right are the top 1 to 5 matched results. The IDs 13, 18, 21, 41 are correctly matched, while the IDs 42 and 47 are incorrect.
IV Experiments
We evaluate the performance of this work from two aspects. First, we evaluate the passenger ReID on our own dataset and compare it with other state-of-the-art approaches. The training is on a single RTX4090 24GB GPU with the batch size, epochs, and initial learning rate set to 8, 60, and 0.0001 with a decay rate 0.1 after 30 epochs, respectively, while retaining some default settings for PFD [34] and TransReID [51]. Next, we run the simulation for transit OD data collection along a manually configured bus route on an NVIDIA Jetson AGX Orin. The confidence threshold of HSDM is set to 0.12 based on the confidence distribution of the training data. This value maximizes the discrimination between positive and negative samples. The logarithmic transformation parameter $k=20$. The weights $\sigma_{1},\sigma_{2},\sigma_{3}$ of the joint similarity are set to $0.2,0.6,0.2$, respectively. To speed up processing, only 8 images are randomly sampled to form a passenger’s tracklet. The loss function follows the same setting in [48]. Performance metrics include Rank-1, 5, 10, 20, mean Average Precision (mAP), and the Cumulative Matching Characteristic (CMC) curve.
IV-A Bus ReID Dataset
To address the current shortage of datasets for occluded re-identification, particularly within public transportation systems, we introduce a new dataset constructed from real-time footage captured over several hours from surveillance cameras mounted at the front and rear doors of buses. The dataset contains over 17,000 images of 157 passengers, with 97 passengers in the training set and 60 in the validation set. It focuses specifically on various types and degrees of occlusions, including those between passengers and between passengers and objects. The dataset also includes images from multiple viewpoints and varying lighting conditions, posing a diverse set of challenges for model training and improving the model’s generalization capabilities in complex environments.
IV-B Evaluation on ReID Performance
The performance of TransitReID on traditional ReID tasks is presented in Table I, where it is compared with several state-of-the-art methods, including PGFA [18], HOReID [52], GRL [48], TransReID [51], PFD [34], and CLIP-ReID [53]. TransitReID achieves the highest accuracy among all comparative methods, improving Rank-1 accuracy by 4.7% and mAP by 8.9%. It is noteworthy that PGFA, HOReID, TransReID, and PFD are specifically designed for occluded ReID, which explains their superior performance compared to the vision-language model CLIP-ReID. Although our baseline model GRL is not specifically designed for occluded cases, it still outperforms some occluded methods due to its video-based ReID approach that learns temporal correlations. Our method not only incorporates temporal continuity information but also selectively focuses on less occluded regions from clear viewpoints to extract useful information. The experimental results show the effectiveness of our method in handling multi-view and complex occlusion scenarios specific to passenger ReID in transit environments.
TABLE I: Comparison with related methods.
Matching Rate (%)
Rank 1 5 10 20 mAP
TransitReID (Ours) 88.3 96.7 100 100 92.5
PGFA[18] 60.9 70.2 75.3 80.4 44.3
HORelD[52] 69.0 75.6 78.2 82.0 58.8
GRL[48] 75.0 96.7 100 100 83.6
TransRelD[51] 83.6 87.1 89.1 91.4 78.8
PFD[34] 78.8 84.0 86.6 89.3 74.6
CLIP-ReID[53] 55.4 68.4 75.0 82.8 45.7
Figure 11: Cumulative Matching Characteristic curve of different ReID methods.
Fig. 11 presents the CMC curves of TransitReID and its comparative methods from Rank-1 to Rank-20. Our method achieves nearly 100% accuracy after Rank-9. Notably, the baseline model GRL attains almost identical accuracy with ours after Rank-5, demonstrating its high performance in video-based ReID tasks. However, TransitReID significantly outperforms GRL from Rank-1 to Rank-4, validating its effectiveness in suppressing occlusion interference. From another perspective, it enhances the baseline model’s matching capability on ‘hard cases’, where the target gallery person often appears within the top 5 matches but fails to achieve the best match due to difficulties in extracting key features and confusion from very similar samples.
Quality Information Extraction
TransitReID employs a grading model to assess images under different occlusion and viewpoint conditions. The assigned scores reflect image quality, where higher-quality images receive greater weights, contributing more information to compensate for the information loss in low-quality images. Fig. 12 visualizes the scores of four tracklet groups. Each image is divided into three segments to represent different body parts; however, this division may not precisely align with the actual body positions and is used solely for visualization purposes.
The first two rows and the last two rows correspond to eight consecutive images of two passengers, but the second and fourth rows include manually added occlusions of random size and position to better illustrate score variations. The manually occluded regions receive lower scores compared to the original images. In the third row, the passenger’s leg region receives lower scores due to a partially obstructing umbrella held in hand, whereas the exposed head and torso receive higher scores.
In the first row, from the fourth image onward, the score for the head region becomes higher than that for the body and legs, even though there is no external occlusion. This can be attributed to the increasing angle relative to the camera as the passenger fully boards the bus. This change creates a ‘folding’ effect in the image, compressing the available visual information, which can be considered a form of ‘self-occlusion’. This demonstrates that the proposed method effectively handles both external occlusions and varying viewpoints.
Figure 12: Score visualization after applying the grading model. Each image is divided into three segments, corresponding to three body parts. The score is mapped to colors from dark blue to red, where a higher score indicates higher quality.
TABLE II: Performance across different numbers of bus stops without / with HSDM.
Stop ID 5 6 7 8 9 10 11 12 13 14 15
Rank-1 0.87 / 0.90 0.87 / 0.90 0.87 / 0.91 0.88 / 0.91 0.86 / 0.90 0.88 / 0.92 0.89 / 0.92 0.88 / 0.91 0.90 / 0.92 0.89 / 0.90 0.89 / 0.90
Rank-5 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
Rank-10 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
IV-C Evaluation on Dynamic Transit OD Data Collection
We randomly assign the 60 passengers in the validation set to several bus stops and randomly determine their boarding and alighting behaviors. Fig. 13 illustrates a simulation with 10 bus stops, where no passengers are on the bus before the starting stop, and passengers can only board at this stop. Similarly, at the terminal stop, all passengers on the bus are required to alight. At each stop, passengers at the stop first board the bus, followed by passengers on the bus alighting. It is worth noting that this setup may differ from real-world scenarios, where boarding and alighting occur simultaneously. As a result, the number of gallery samples at each stop in the simulation is overestimated, increasing the matching difficulty but providing a more rigorous evaluation of our method.
Fig. 9 presents the results of OD data collection. The first and last rows represent the 10 boarding and alighting stops, while the middle row represents the 60 passengers. Prediction occurs between passengers and the alighting stops, where gray lines indicate correct matches and green lines indicate incorrect matches. An incorrect match means a passenger is incorrectly matched with an alighting passenger, leading to a false prediction of the alighting stop. Out of the 60 matches, only 3 are incorrect, resulting in a total accuracy of 95%. It is important to note that this accuracy differs from the accuracy in the Ablation Study. The former represents the OD collection accuracy, while the latter represents the ReID retrieval accuracy.
Some retrieval results are shown in Fig. 10. The examples illustrate matching performance under various conditions, including different viewpoints, passenger occlusions, and both daytime and nighttime scenarios. The boarding and alighting images of the target passengers are highlighted with red and green boxes, respectively. From left to right are the top-1 to top-5 retrieval results.
The IDs 13, 18, 21, and 41 are correctly matched cases, where the top-1 retrieval result corresponds to the ground truth marked by the green box. This demonstrates the effectiveness of our method in extracting discriminative features under diverse and complex environments.
Two failure cases are also presented, corresponding to IDs 42 and 47. Specifically, ID 47 alights earlier than ID 42, and the ground truth of ID 42 is mistakenly matched as the top-1 result of ID 47. As a result, when ID 42 actually alights, its ground truth is no longer available for matching. This indicates that while the proposed HSDM strategy can reduce incorrect matches to some extent, it cannot achieve 100% accuracy. The choice of the confidence threshold for matching remains a critical factor. Nevertheless, we argue that an accuracy exceeding 90% in the simulation experiment is sufficient to validate the effectiveness of our approach in transit data collection applications.
Figure 13: Boarding and alighting simulation with 60 passengers in 10 bus stops.
IV-D Inference Time Analysis on Edge Devices
We conduct experiments on an NVIDIA Jetson AGX Orin to simulate the inference time of an edge computing device in a realistic transit scenario. The results are presented in Table III. Board/Alight indicates the number of passengers boarding or alighting at each stop. $T_{F}$ represents the total inference time for feature extraction, and $T_{M}$ represents the total time for ReID retrieval. The last row shows the average inference time per passenger. The processing time per passenger is approximately 3 seconds, mainly constrained by the computational performance of the edge device. Nevertheless, We implement multi-threaded programming to enable real-time operation in practical smart transit applications.
As shown in Fig. 14, the video thread, connected to the transit surveillance system, captures video frames from the front and rear doors and feeds them into two separate detection threads. In these detection threads, real-time processing is achieved due to the high efficiency of the YOLO model. The featurize thread receives the detection masks, extracts feature embeddings, and handles storage, matching, and HSDM execution. The visualization thread is designed for debugging purposes and is not utilized in the practical deployment.
The real-time performance of the detection threads ensures smooth operation within the transit system. Meanwhile, the featurize thread does not require strict real-time execution—as long as the processing is completed before the next stop. For example, if 10 passengers board the bus at stop $i$, the entire processing should be completed within 30 seconds, before the bus arrives at stop $i+1$.
Figure 14: Multi-threaded programming.
TABLE III: Statistics of each stop’s information and processing time.
Stop ID 1 2 3 4 5
Board/Alight 8/0 8/1 3/1 7/6 9/2
$T_{F}$ 33.4 30.3 13.2 42.8 36.0
$T_{M}$ 1.55e-3 1.38e-3 6.30e-3 2.81e-3
$T_{mean/P}$ 3.37 3.30 3.31 3.00
Stop ID 6 7 8 9 10
Board/Alight 8/2 5/11 7/9 5/7 0/21
$T_{F}$ 32.7 52.8 52.0 39.3 68.2
$T_{M}$ 2.72e-3 1.22e-2 1.00e-2 7.92e-3 2.14e-2
$T_{mean/P}$ 3.27 3.30 3.06 2.81 3.25
IV-E Ablation Study
The effectiveness of the proposed HSDM mechanism is evaluated through simulations with the number of bus stops $S$ ranging from 5 to 15. Each configuration involves 20 repeated experiments, and the average matching rate across all stops is calculated. The results are shown in Table II.
Without HSDM, the Rank-1 matching rate increases very slightly as the number of bus stops increases. This is attributed to the reduction in gallery samples at each stop, which enhances the matching success rate. With HSDM, the Rank-1 rate increases across all settings, demonstrating the mechanism’s capability to handle low-confidence matches by providing these samples with additional opportunities to achieve the global optimum.
Notably, with HSDM, the Rank-1 rate remains stable regardless of the number of bus stops. This stability is achieved through the cold storage mechanism, which ensures that gallery samples at each stop are influenced not only by the number of passengers on board but also by their matching confidence. Consequently, the number of matching samples are larger than the actual number of passengers onboard, further validating the robustness of the proposed method under varying numbers of bus stops.
For Rank-5 and Rank-10, no significant changes are observed with or without HSDM, indicating the high accuracy of our ReID algorithm for simple samples.
V Conclusion
This paper presented TransitReID, a novel framework for individual-level and dynamic transit OD data collection. The proposed solution addressed two critical challenges in transit environments: (1) severe occlusion through an innovative region-attention ReID mechanism that integrates a VAE-based grading model to dynamically weight visible body regions, coupled with adaptive feature fusion to preserve discriminative information; (2) continuous operation requirements via a Hierarchical Storage and Dynamic Matching (HSDM) architecture that incorporated Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. Furthermore, a multi-threaded design is implemented to achieve near real-time operation on transit vehicles equipped with edge devices. As all features are processed and stored locally, passenger privacy is effectively protected. Comprehensive evaluations demonstrated the framework’s effectiveness, achieving state-of-the-art performance with 88.3% Rank-1 accuracy on our proposed transit ReID dataset and maintaining on average 90% OD estimation accuracy in real-world transit simulations.
While real video data are now available, the configuration of bus stops remains a manual process. Future work will involve deploying this system in operational public transport vehicles to conduct further tests and refinements, ensuring real-time performance. Moreover, there is a pressing need to develop lighter algorithms based on this research to facilitate real-time execution on edge devices.
VI Acknowledgments
The authors would like to express their sincere gratitude to the Urban Transportation Associates (UTA) for the valuable insights and support in this research study.

### v2

@@
 Abstract
-Transit Origin-Destination (OD) data are essential for transit planning, particularly in route optimization and demand-responsive paratransit systems.
-Traditional methods, such as manual surveys, are costly and inefficient, while Bluetooth and WiFi-based approaches require passengers to carry specific devices, limiting data coverage.
-On the other hand, most transit vehicles are equipped with onboard cameras for surveillance, offering an opportunity to repurpose them for edge-based OD data collection through visual person re-identification (ReID).
-However, such approaches face significant challenges, including severe occlusion and viewpoint variations in transit environments, which greatly reduce matching accuracy and hinder their adoption.
-Moreover, designing effective algorithms that can operate efficiently on edge devices remains an open challenge.
-To address these challenges, we propose TransitReID, a novel framework for individual-level transit OD data collection.
-TransitReID consists of two key components: (1) An occlusion-robust ReID algorithm featuring a variational autoencoder guided region-attention mechanism that adaptively focuses on visible body regions through reconstruction loss-optimized weight allocation; and (2) a Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient and robust transit OD matching which balances storage, speed, and accuracy.
-Additionally, a multi-threaded design supports near real-time operation on edge devices, which also ensuring privacy protection.
-We also introduce a ReID dataset tailored for complex bus environments to address the lack of relevant training data.
-Experimental results demonstrate that TransitReID achieves state-of-the-art performance in ReID tasks, with an accuracy of approximately 90% in bus route simulations.
+Transit Origin-Destination (OD) data are fundamental for optimizing public transit services, yet current collection methods, such as manual surveys, Bluetooth/WiFi tracking, or Automated Passenger Counters, are either costly, device-dependent, or incapable of individual-level matching.
+Meanwhile, onboard surveillance cameras already deployed on most transit vehicles provide an underutilized opportunity for automated OD data collection.
+Leveraging this, we present TransitReID, a novel framework for individual-level and occlusion-resistant passenger re-identification (ReID) tailored to transit environments.
+Our approach introduces three key innovations: (1) an occlusion-robust ReID algorithm that integrates a variational autoencoder–guided region-attention mechanism and selective quality feature averaging to dynamically emphasize visible and discriminative body regions under severe occlusions and viewpoint variations; (2) a Hierarchical Storage and Dynamic Matching (HSDM) mechanism that transforms static gallery matching into a dynamic process, balancing accuracy, memory, and speed in real-world bus operations; and (3) a multi-threaded edge implementation that enables near real-time OD estimation while ensuring privacy by processing all data locally.
+To support research in this domain, we also construct a new Transit ReID dataset with over 17,000 images captured from bus front/rear cameras under diverse occlusion and viewpoint conditions.
+Experimental results demonstrate that TransitReID achieves state-of-the-art performance, with R-1 accuracy of 88.3% and mAP of 92.5%, and further sustains 90% OD estimation accuracy in bus route simulations on NVIDIA Jetson edge devices.
+This work advances both the algorithmic and system-level foundations of automated transit OD collection, paving the way for scalable, privacy-preserving deployment in intelligent transportation systems.
 Index Terms:
@@
 It aids in the analysis of relationships and impacts among regional economic development, infrastructure, and urban mobility [1].
-This data represents the demand of passengers and is needed to design the bus routes efficiently to minimize travel time, reduce operating costs, and driver scheduling.
+For public transportation systems, OD data represents passenger demand and plays a crucial role in designing routes more efficiently to minimize travel time, reduce operating costs, and optimize driver scheduling.
 As shown in Fig.
@@
 More advanced methods, such as those utilizing mobile phone data[3] and Bluetooth technology [4], require passengers to carry specific devices with WiFi or Bluetooth enabled, which limits coverage and might raise privacy concerns due to the unique identifiers of the device.
-Automated Passenger Counters (APCs) [5] also assist in estimating OD data, however, they can only capture the aggregate number and locations of boarding and alighting passengers.
-They fail to match passengers at the individual level, which is crucial for studies focused on individual passenger behavior.
-Meanwhile, most transit vehicles in the US are already equipped with onboard cameras for liability and surveillance purposes, presenting a valuable opportunity to repurpose them as smart sensors for transit OD data collection.
-By integrating these cameras with person re-identification (ReID) techniques and edge computing devices, seamless individual-level matching can be achieved [6].
-By integrating sensors such as GPS, the matching results can be readily translated into transit OD data.
-Despite the advantages of camera-based methods over prevailing techniques, their adoption in transit OD data collection remains limited due to challenges posed by the transit vehicle environment, particularly occlusion issues.
-For example, during peak times, boarding and alighting passengers may obscure each other from the camera view, leading to information loss and background confusion.
-Recent research increasingly focuses on occlusion issues in ReID, broadly categorized into two types: methods based on local attention mechanisms and those based on feature recovery.
+Automated Passenger Counters (APCs) [5] also assist in estimating OD data, however, they can only capture the counts and locations of boarding and alighting passengers and fail to match passengers at the individual level, which is crucial for understanding detailed travel behavior.
+From a transit planner’s perspective, individual-level OD information provides significant advantages over aggregated OD data.
+It enables more accurate demand forecasting by revealing detailed movement patterns, supports the design of tailored services for different passenger groups such as commuters, students, and occasional travelers, and facilitates the evaluation of OD-related performance indicators across user segments.
+Moreover, it helps identify irregular behaviors, such as fare evasion or abnormal travel trajectories, and can assist law enforcement in locating suspects or missing persons, thereby improving both operational efficiency and public safety.
+Figure 1: Traditional methods for estimating transit OD data and our method.
+To address these needs, video-based re-identification (ReID) techniques offer a promising solution.
+Unlike device-dependent approaches, video-based ReID captures continuous passenger trajectories without requiring user cooperation and, when combined with anonymization strategies, can achieve this while mitigating privacy concerns.
+Integrating ReID into OD estimation frameworks can thus significantly enhance the accuracy, coverage, and utility of OD data, enabling transit planners to design smarter and more responsive public transportation systems.
+Meanwhile, most transit vehicles in the US are already equipped with onboard cameras for liability and surveillance purposes, presenting a valuable opportunity to repurpose them as smart sensors for video-based transit OD data collection [6].
+By integrating these cameras with ReID techniques, edge computing devices, and GPS, seamless individual-level matching and OD data collection can be achieved.
+Despite the advantages of camera-based methods over existing techniques, their adoption in transit OD data collection remains limited due to several challenges, particularly occlusion.
+For example, during peak times, passengers may obscure each other from the camera view, leading to information loss and background confusion.
+Recent research on ReID has increasingly emphasized occlusion, with existing approaches broadly falling into two categories: local attention–based methods and feature recovery–based methods.
 The former constructs attention mechanisms by focusing on different local areas to suppress interference from occluded parts [7], while the latter attempts to recover information from occluded areas by associating temporal or spatial information [8, 9].
-Fixed local regions are used to differentiate various body parts of pedestrians and assign scores to different areas using a quality predictor in [10].
-However, this fixed-area approach is not suitable for passenger matching tasks on transit vehicles under dynamic viewpoints.
-As the distance and angles between passengers and cameras change over time, the regions of their body parts shift throughout the video.
-Figure 1: Traditional methods for estimating transit OD data and our method.
-Due to limitations in power consumption, memory, and computational resources, edge devices are generally incapable of executing advanced and complex analytical algorithms[11, 12].
-Fast and accurate estimation of OD (origin-destination) data is essential for enabling real-time passenger flow analysis in transit systems.
-However, most deep learning algorithms have relatively long inference times on edge devices, leading to latency and performance degradation.
-Some efforts have focused on model lightweighting techniques, such as knowledge distillation and pruning, or on optimizing neural network architectures specifically for edge hardware[13].
-Nevertheless, lightweight models often suffer from a trade-off in performance, with reduced accuracy to some extent.
-What’s more, unlike traditional ReID studies, our work focuses on the problem of transit OD data collection, hence the gallery, shown in Fig.
+Although some methods achieve strong performance on general ReID benchmarks, they still struggle with dynamic viewpoints and complex occlusion patterns, and often fail to transfer effectively to real-world application scenarios such as transit vehicles.
+Additionally, unlike traditional ReID studies, our work focuses on the problem of transit OD data collection, hence the gallery, shown in Fig.
 2, is dynamic, changing with the variations in passengers boarding and alighting at each station.
-Consequently, ReID evolves from a problem of matching 1 query to $N$ options, into a dynamic 1 to $N_{1},N_{2},...N_{K}$ problem, where $N_{k}$ represents the size of the gallery updated at station $k$.
+Consequently, transit ReID evolves from a problem of matching 1 query to $N$ options, into a dynamic 1 to $(N_{1},N_{2},...,N_{K})$ problem, where $N_{k}$ represents the size of the gallery updated at station $k\in K$.
 To address this, we design a hierarchical storage and dynamic matching (HSDM) method to balance the trade-offs among matching accuracy, storage pressure, and processing speed.
-To jointly tackle the aforementioned issues of occlusion and the constraints of online adaptive matching, we propose TransitReID, an individual-level and robust transit OD data collection method that incorporates occlusion-resistant and dynamic passenger ReID.
+The limited computational resources of edge devices present an additional challenge in executing advanced and complex analytical algorithms [10, 11].
+Fast and accurate estimation of OD data is crucial for minimizing latency in passenger flow analysis.
+To address this, some efforts have focused on model lightweighting techniques, such as knowledge distillation and pruning, or on optimizing neural network architectures specifically for edge hardware [12].
+Nevertheless, lightweight models often face a trade-off between efficiency and accuracy, typically leading to a certain degree of performance degradation.
+Figure 2: Segmentation results after detection and tracking.
+To jointly tackle the aforementioned issues, we propose TransitReID, an individual-level and robust transit OD data collection method that incorporates occlusion-resistant and dynamic passenger ReID.
 The overall framework is shown in Fig.
-3, featuring two principal technical innovations: (1) An occlusion-robust ReID algorithm that employs a region-attention module to adaptively weight less obscured body segments with clearer viewpoints, guided by a co-trained Variational Autoencoder (VAE) grading model.
+3, featuring two principal technical innovations: (1) An occlusion-robust ReID algorithm employing a region-attention module to adaptively weight less-obscured body segments with clearer viewpoints, guided by a co-trained Variational Autoencoder (VAE) grading model robust to occlusion and viewpoint variations.
 The VAE’s reconstruction loss directly informs attention weight allocation, emphasizing visually coherent regions.
-This is followed by a feature fusion strategy that preserves discriminative information through concatenated weighted averages of region embeddings.
+This is followed by a Selective Quality Feature Averaging (SQFA) strategy that preserves discriminative information through concatenated weighted averages of region embeddings.
 (2) A Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient transit OD matching, which incorporates Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed.
@@
 This variability is crucial for training a robust and generalizable model under transit bus operational scenarios.
-Therefore, we construct a passenger ReID dataset collected from real bus operational settings, including footage from cameras at the front and rear doors.
-This dataset captures 17,636 images of 157 passengers.
-In essence, the major contributions of this work are as follows:
-  • Onboard transit surveillance systems are combined with computer vision technology to collect transit OD data, achieving individual-level precision in the construction of the OD matrix.
-  • An attention-based passenger ReID algorithm is specifically designed for transit vehicles scenarios with variable occlusion and viewpoints.
+Therefore, we construct a transit ReID dataset (detailed in Section IV.
+A.) collected from real bus operational settings, including 17,636 images of 157 passengers captured by cameras at the front and rear doors.
+The major contributions of this work are as follows:
+  • We propose TransitReID to enable individual-level transit OD data collection, employing an attention-based passenger ReID algorithm that is robust to variable occlusions and viewpoint changes.
   • A hierarchical storage and dynamic matching (HSDM) mechanism tailored for actual transit scenarios is designed, transforming the traditional static ReID problem into a dynamic ReID challenge.
-  • An efficient programming approach is designed to enable real-time operation on edge devices for smart transit applications.
-In addition, edge-based processing inherently supports privacy protection.
-Figure 2: Segmentation results after detection and tracking.
+  • Introduce an efficient programming approach to enable real-time operation on edge devices for smart transit applications. Edge-based processing inherently supports privacy protection.
+Figure 3: The framework of the proposed TransitReID.
+The surveillance system captures image sequences of the passengers boarding and alighting at different stops.
+After behavior detection, each image is processed by a human parsing network, which divides the image into three body parts: ‘head’, ‘torso’, and ‘legs’.
+Each part is then fed into a grading model and a feature generation model to obtain a quality score and an initial feature representation.
+Finally, a selective quality feature averaging (SQFA) unit integrates these outputs into the final feature vector.
+At boarding, the feature is first stored in the gallery; at alighting, the system employs the hierarchical storage and dynamic matching (HSDM) mechanism to identify the corresponding feature from the gallery.
 II Related Work
 II-A Transit OD Data Collection
-Transit OD data collection refers to the systematic process of gathering information about the starting points (origins) and endpoints (destinations) of individual trips within a public transportation network.
-The most traditional and ubiquitous methods for estimating transit OD data is by surveys [22, 23], which are labor-intensive and have limited response rates.
+Transit OD data collection refers to the systematic process of gathering information on the origins and destinations of individual trips within a public transportation network.
+OD trip matrices can be estimated either from direct trip observations or derived from traffic counts.
+The latter approach, known as OD synthesis, estimates the OD matrix through linear combinations of observed flows [21].
+Common methods include gravity-based models [22], entropy models [23, 24], and statistical approaches based on least squares [25] or maximum likelihood estimation [26].
+However, these methods are typically applied to vehicular traffic.
+The most ubiquitous methods for estimating passenger-level transit OD data is by surveys [27, 28], which are labor-intensive and have limited response rates.
 In [29], an integrated program based on Automated Fare Collection (AFC) systems was developed that can generate OD matrices and passenger load profiles by making some assumptions about round trips.
@@
 RFID technology was utilized to create a Boarding-In and Boarding-Out (BIBO) system for passenger monitoring in [32].
-Moreover, mobile phone data [3], Bluetooth technology [4], and multi-source approaches – such as combining GIS data with passenger IC card data to construct OD matrices, as demonstrated in [28] – have been utilized.
-However, these methods face challenges, including high costs, significant maintenance requirements [2], and the need for passengers to carry specific devices, which limits data coverage.
-Additionally, the extensive use of these systems in public transport can lead to privacy concerns [29], such as exposing details about ticket types, passenger categories, travel behaviors, and purchasing methods.
-While most methods capture only the number and locations of boarding and alighting passengers, they lack the precision to individually track and match passengers, crucial for detailed behavioral studies.
-Video-based passenger re-identification (ReID) technology combined with transit surveillance cameras provides a potential solution to this problem.
-In [6], the YOLO algorithm is utilized to detect passengers, generating feature vectors through a neural network and selecting matches based on vector distance, but this work ignores the passive influence of occlusion.
-Several works [30, 31, 32] address the problem by only collecting information from head images of passengers inside the bus.
-However, this approach overlooks the potential rich information from other regions, lacks discrimination power against similar headgear, and is not robust to changes in lighting.
+Moreover, mobile phone data [3], Bluetooth technology [4], and multi-source approaches — such as integrating GIS data with passenger IC card data to construct OD matrices, as demonstrated in [33] — have been employed.
+However, these methods face several challenges, including high costs, substantial maintenance requirements [2], and the dependence on passengers carrying specific devices, which limits data coverage.
+In addition, widespread deployment of such systems in public transportation raises privacy concerns [34], potentially revealing sensitive information.
+While most existing methods capture only aggregate counts and locations, they lack the precision required to track and match individual passengers, which is essential for detailed behavioral analysis.
+Video-based passenger re-identification using transit surveillance cameras offers a promising solution.
+In [6], the YOLO algorithm was employed to detect passengers, followed by feature extraction through a neural network and matching based on vector distances.
+However, this study does not consider the passive effects of occlusion.
+Other works [35, 36, 37] addressed this issue by focusing solely on head images of passengers inside the bus.
+Yet, they neglects potentially informative regions of the body, exhibits limited discriminative power against similar headgear, and is sensitive to variations in lighting conditions.
 Our proposed methodology not only maximizes the useful information extracted from each passenger through an attention mechanism but also enhances privacy protection by processing data locally on onboard edge devices, ensuring that only high-level features are stored.
 II-B Occluded Re-Identification
-Occlusion is by far the major problem in the computer vision-based person ReID task.
-Most of the occluded person re-identification methods rely on convolutional neural networks (CNNs)[7].
-These aim at body part-level feature measurement, for example, a region-based quality predictor was proposed and trained with the regional feature generation model to get the weighted feature representation in [10].
-However, the regions defined in this work is fixed, thus not suitable for dynamic perspectives.
-A pose-guided feature alignment (PGFA)[18] utilized human pose landmarks to build an attention map, guiding the model to extract useful information from non-occluded parts.
-Similarly, [33, 34, 35, 36] utilize posture-guided attention mechanisms to discern visible areas.
+Occlusion is one of the most significant challenges in vision-based person ReID.
+Some of the occluded person re-identification methods rely on convolutional neural networks (CNNs)[7].
+They aim at body part-level feature measurement, for example, a region-based quality predictor was proposed and trained with the regional feature generation model to get the weighted feature representation in [38].
+However, the regions defined in this work is fixed, thus not suitable for passenger matching tasks on transit vehicles under dynamic viewpoints.
+As the distance and angles between passengers and cameras change over time, the regions of their body parts shift throughout the video.
+A pose-guided feature alignment method (PGFA)[17] utilized human pose landmarks to build an attention map, guiding the model to extract useful information from non-occluded parts.
+Similarly, [39, 40, 41, 42] utilized posture-guided attention mechanisms to discern visible areas.
 These methods partition the ReID problem into two segments: initially detecting key areas, then selectively extracting regional features for matching.
@@
 Additionally, the lack of scenario-specific training data further limits the models’ generalizability.
-Moreover, they address static ReID problems, whereas the gallery in a transit vehicle operating environment is inherently dynamic.
+Moreover, they address static ReID problems, whereas the gallery in a transit vehicle is inherently dynamic.
 We design a hierarchical storage and dynamic matching (HSDM) mechanism specifically adapted for dynamic ReID, detailed in Section III.
 F.
-Figure 3: The Framework of the Proposed TransitReID.
-In Dynamic Transit ReID, the surveillance system captures image sequences of the same passenger boarding and alighting at different stops.
-These images are processed by a feature generation network to extract one-dimensional feature representations.
-If the passenger is boarding, the features are temporarily stored in the gallery.
-If the passenger is alighting, the system retrieves a matching result from the gallery through the hierarchical storage and dynamic matching (HSDM) mechanism.
-In the Feature Generation Network, the input consists of an image sequence of a passenger.
-After action detection and segmentation, each mask is processed by a human parsing network, which divides the image into three body parts: ‘head’, ‘torso’, and ‘legs’.
-Each part is then fed into a grading model and a baseline model to obtain a quality score and an initial feature representation.
-Finally, a feature averaging unit integrates these outputs into the final 3*2048 length feature vector.
 III Methodology
@@
 In the following sections, we focus on the transit bus environment, which is equipped with two cameras: a front-door camera $C_{f}$ and a rear-door camera $C_{r}$, which capture sequences of images over a time interval $T$.
-The image sets captured by each camera are:
-$I_{f}=\{i_{f,1},i_{f,2},...,i_{f,T}\},\quad I_{r}=\{i_{r,1},i_{r,2},...,i_{r,T}\}.$
-Each image may contain multiple passengers. The gallery, $G$ consists of all passenger images extracted from the raw images recorded by both cameras while onboard the bus:
-$G=G_{f}\cup G_{r}=\{g_{f,1},g_{f,2},...,g_{f,M_{f}}\}\cup\{g_{r,1},g_{r,2},...,g_{r,M_{r}}\},$
+Each image may contain multiple passengers.
+The gallery $G$ consists of all passenger images extracted from the raw images recorded by both cameras while onboard the bus:
+$G=G_{f}\cup G_{r}=\{g_{f,1},g_{f,2},...,g_{f,M_{f}}\}\cup\{g_{r,1},g_{r,2},...,g_{r,M_{r}}\}$ (1)
 where $M_{f}$ and $M_{r}$ are the number of on board passengers captured by the front and rear cameras, respectively, and $M=M_{f}+M_{r}$ is the total gallery size.
-At each bus stop, the query set, $Q$, consists of passengers departing from either the front or rear door:
-$Q=Q_{f}\cup Q_{r}=\{q_{f,1},q_{f,2},...,q_{f,N_{f}}\}\cup\{q_{r,1},q_{r,2},...,q_{r,N_{r}}\},$
-where $N_{f}$ and $N_{r}$ are the number of passengers alighting from the front and rear doors, respectively, and $N=N_{f}+N_{r}$ is the total number of alighting passengers.
-Each passenger $p_{k}$ is mapped to a high-dimensional feature space via a function:
-$\phi:p_{k}\rightarrow\mathbb{R}^{d},$
+At each bus stop, the query $Q$ updates passengers departing from either the front or rear door:
+$Q=Q_{f}\cup Q_{r}=\{q_{f,1},q_{f,2},...,q_{f,N_{f}}\}\cup\{q_{r,1},q_{r,2},...,q_{r,N_{r}}\}$ (2)
+where $N_{f}$ and $N_{r}$ are the number of passengers alighting from the front and rear doors, respectively, and $N=N_{f}+N_{r}$ is the total query size.
+Each passenger $p$ is mapped to a high-dimensional feature space via a function:
+$\phi:p\rightarrow\mathbb{R}^{d}$ (3)
 where $d$ is the feature space dimension.
-For each alighting passenger $q_{k}\in Q$, we determine the best match from the gallery $G$ by computing the distance between $\phi(q_{k})$ and $\phi(g_{j})$ for all $g_{j}\in G$, selecting the closest match:
-$\mathbf{min}\ distance(\phi(q_{k}),\phi(g_{j})),\quad\forall g_{j}\in G.$
+For each alighting passenger $q_{n}\in Q$, we determine the best match from the gallery $G$ by computing the distance between $\phi(q_{n})$ and $\phi(g_{m})$ for all $g_{m}\in G$, selecting the closest match:
+$\hat{g}=\arg\min_{g_{m}\in G}\,distance\big{(}\phi(q_{n}),\phi(g_{m})\big{)}$ (4)
 III-B Detection and Tracking
@@
 $\big{(}S\in R_{\text{door}}\big{)}\land\big{(}E\in R_{\text{door}}\big{)}$
-These equations correspond to trajectories of alighting, boarding, moving while inside the bus, and moving while remaining outside.
+These equations correspond to trajectories of 1) alighting, 2) boarding, 3) moving inside the bus, and 4) moving while remaining outside.
+Taking the second formula as an example, a passenger is classified as boarding only if they first appear in $R_{door}$, then leave from the $R_{inside}$ region, and do not return to $R_{door}$.
+This processing effectively filters out false boarding and alighting behaviors, such as when a passenger lingers near the door but ultimately does not board the bus.
 Only trajectories satisfying the first two cases are retained, with their corresponding locations recorded as key-value pairs.
@@
 III-D Grading Model
-Passengers’ local features should be weighted differently to account for various occlusion and viewpoint scenarios.
-Existing work [10] utilizes a convolutional network including a convolution layer and a fully connected layer to compute local scores.
-[47] designs a spatial-patch contrastive (SPC) loss to compute aggregation maps to focus on different areas.
-Our grading model uses a Variational Autoencoder (VAE) to assign scores to images of different body parts, where a higher score indicates better image quality, with less occlusion and a clearer viewpoint.
+We design a grading model using a Variational Autoencoder (VAE) to assign scores to images of different body parts, where a higher score indicates better image quality, with less occlusion and a clearer viewpoint.
 This score will be used to calculate image weights in the subsequent processing stages.
@@
 5.
-The image input passes through an encoder, which consists of a convolutional layer, followed by a max-pooling layer after each convolution.
-The features are then processed by residual modules to obtain a flattened feature map.
-Two fully connected layers calculate the mean and variance of the latent space, introducing randomness to the model and enabling better learning of reconstruction features.
-The decoder consists of a series of transposed convolutional layers and residual modules which output the reconstructed image through a sigmoid function.
-During the training, the grading model is jointly trained with the feature extraction network Global-guided Reciprocal Learning (GRL) [48].
+During the training, the grading model is jointly trained with the feature extraction network.
 The underlying logic is straightforward: through training, the VAE is enabled to better reconstruct images of higher quality and poorly reconstruct those with severe occlusions and bad viewpoints.
-The image score, $I_{S}$, is calculated based on the joint similarity between the reconstructed and original images:
-$I_{S}=\sigma_{1}SSIM+\sigma_{2}HS+\sigma_{3}MSE$ (2)
+Note that this VAE structure is inherently robust to complex occlusion patterns and dynamic viewpoints, as it relies on the holistic integrity of the image, regardless of the viewing angles or how the occlusion occurs.
+The image score $I_{S}$ is calculated based on the joint similarity between the reconstructed and original images:
+$I_{S}=\sigma_{1}\text{SSIM}+\sigma_{2}\text{HS}+\sigma_{3}\text{MSE}$ (6)
 To comprehensively assess the image reconstruction effect, the above formula considers three measurement factors.
-The Structural Similarity Index Measure (SSIM) [49] is an indicator to measure the visual similarity of two images based on three comparison parameters: luminance, contrast, and structure.
-Histogram Similarity (HS) focuses on the pixel value distribution of images, calculating the similarity of the histograms of the raw and reconstructed images.
-The Mean Squared Error (MSE) reflects the pixel value differences between the two images.
-The score is then fed into the Score Mapping for further process.
+The Structural Similarity Index Measure (SSIM) [53] measures the visual similarity of two images.
+The Histogram Similarity (HS) calculates the similarity of the histograms.
+The Mean Squared Error (MSE) reflects the pixel value differences.
 Score Mapping
-A Logarithmic Transformation is used to map the original scores to a range from 0 to 1.
-Additionally, the logarithmic transformation compresses a wide score range into a smaller output range, which benefits our scenarios, where severe occlusions are of greater concern, the mapped scores decrease rapidly, helping to effectively suppress high-occlusion situations.
-For cases with minor occlusions, we aim to preserve the features of these images to the greatest extent.
-Equation (3) defines a mapping $\mathcal{L}:x\rightarrow y$, where $x$ is the input score and $y$ is the transformed score.
-$min/max$ is the upper/lower bound of the input.
+A logarithmic transformation is applied to map the original scores to a range of 0 to 1.
+It compresses a wide score range into a narrower output range, which is particularly beneficial in scenarios where severe occlusions are of greater concern.
+In such cases, the lower scores decrease rapidly, effectively suppressing the impact of high-occlusion situations.
+Conversely, for images with minor occlusions, the transformation preserves their features to the greatest extent.
+Equation (3) defines a mapping $\mathcal{L}:x\mapsto y$, where $x$ is the score and $\min/\max$ is the upper/lower bound of the input.
 Fig.
 6 displays the mapping curves at different values of $k$.
-$y=\frac{\log(1+k\cdot(x-\min))}{\log(1+k\cdot(\max-\min))}$ (3)
+$\mathcal{L}(x)=\frac{\log(1+k\cdot(x-\min))}{\log(1+k\cdot(\max-\min))}$ (7)
 Figure 6: Logarithmic transformation curves with different $k$ values.
@@
 It effectively utilizes both global and local feature correlations within video frames.
-This Global-guided Reciprocal Learning (GRL) framework incorporates two key modules: the Global-guided Correlation Estimation (GCE) and the Temporal Reciprocal Learning (TRL).
-The GCE module generates correlation maps to distinguish between high- and low-correlation features under global guidance, enhancing the model’s focus on relevant features.
-The TRL module optimizes the use of these features over time, improving the discrimination of temporal information.
-The features for three body parts Head, Torso, Legs ($H,T,L$), are processed separately to obtain the feature sets $\mathfrak{F}=\{F_{H},F_{T},F_{L}\}$, where $\underset{k=H,T,L}{F_{k}}=\{f_{1}^{k},f_{2}^{k},...,f_{l}^{k}\}$.
-The score sets output by the grading model are denoted by $\mathcal{S}=\{S_{H},S_{T},S_{L}\}$, where $\underset{k=H,T,L}{S_{k}}=\{s_{1}^{k},s_{2}^{k},...,s_{l}^{k}\}$.
+The features for three body parts Head, Torso, Legs ($H,T,L$), are processed separately to obtain the feature sets:
+$\mathcal{F}=\{F_{k}=\{f_{1}^{k},f_{2}^{k},\dots,f_{l}^{k}\}\}_{k\in\{H,T,L\}}$ (8)
+The score sets output by the grading model are denoted by:
+$\mathcal{S}=\{S_{k}=\{s_{1}^{k},s_{2}^{k},\dots,s_{l}^{k}\}\}_{k\in\{H,T,L\}}$ (9)
 We use $l=8$ frames per person for consistency.
-Figure 7: Selective Quality Feature Averaging.
+Figure 7: Selective quality feature averaging (SQFA).
 Each image is attached with a score reflecting its degree of occlusion.
-Averaging all frames would suppress those of high quality.
-Averaging only frames with scores above a threshold, however, would preserve the features of high quality.
+Averaging only frames with scores above the threshold (as shown in the boxes) would preserve the features of high quality.
 Unlike the baseline model [54], we do not consider all the images in a sequence.
@@
 Ultimately, the feature vector for a passenger is calculated by the following formula:
-$\mathcal{F}=\underset{k=H,T,L}{\mathtt{CONCAT}}\;(\frac{F_{k}\odot S_{k}}{l-d})$ (4)
+$\widetilde{F}=\operatorname{CONCAT}_{k\in\{H,T,L\}}\Biggl{(}\frac{F_{k}\odot S_{k}}{l-d}\Biggr{)}$ (10)
 Where $d$ is the number of images in the sequence whose scores fall below a certain threshold.
 $\odot$ is the Hadamard product.
-$\mathtt{CONCAT}$ indicates that vectors from the three parts are concatenated along the last dimension to form a one-dimensional feature vector.
+$\operatorname{CONCAT}$ indicates that vectors from the three parts are concatenated along the last dimension to form a one-dimensional feature vector.
 These features would then be stored in the Facebook AI Similarity Search (FAISS) library [55] which is an optimized library for large-scale high dimensional nearest neighbor similarity search.
@@
 8.
-Figure 8: Structure of hierarchical storage and dynamic matching.
+Figure 8: Structure of hierarchical storage and dynamic matching (HSDM).
 The HSDM consists of 3 major steps as follows, and the processing logic is summarized in Algorithm 1.
@@
 5:  for each passenger $p_{\text{alight}}$ alighting at station $i$ do
-[…]
+6:   $(\text{topK\_IDs},\text{topK\_dists})=\text{FAISS.search}$
 7:   Calculate $\gamma$
@@
 9:     Hot Storage:
-[…]
-[…]
+10:     $p_{\text{alight}}.\text{matched\_ID}=\text{topK\_IDs}[0]$
+11:     Remove $\text{topK\_IDs}[0]$ from FAISS index
 12:   else
 13:     Cold Storage:
-[…]
+14:     $p_{\text{alight}}.\text{temp\_match}=\text{topK\_IDs}[0]$
 15:     $p_{\text{alight}}.\text{alt\_IDs}=\text{topK\_IDs}[:K]$
@@
 17:for each new passenger $p_{\text{alight2}}$ at station $j>i$ (if exists) do
-[…]
+18:  $(\text{topK\_IDs2},\text{topK\_dists2})=\text{FAISS.search}$
 19:  Calculate $\gamma_{2}$
-[…]
+20:  if $\text{topK\_IDs2}[0]$ is in Cold Storage by another $p_{\text{alight}}$ then
 21:   if $\gamma_{2}>\gamma_{\text{old}}$ then
-[…]
+22:     Remove $\text{topK\_IDs2}[0]$ from $p_{\text{alight}}.\text{temp\_match}$
 23:     if $\gamma_{2}>\sigma$ then
-[…]
-[…]
+24:      $p_{\text{alight2}}.\text{matched\_ID}=\text{topK\_IDs2}[0]$
+25:      Remove $\text{topK\_IDs2}[0]$ from FAISS index
 26:     else
-[…]
+27:      $p_{\text{alight2}}.\text{temp\_match}=\text{topK\_IDs2}[0]$
 28:      $p_{\text{alight2}}.\text{alt\_IDs}=\text{topK\_IDs2}[:K]$
@@
 33:for each $p_{\text{alight}}$ still in Cold Storage and not snatched do
-[…]
+34:  $p_{\text{alight}}.\text{matched\_ID}\leftarrow p_{\text{alight}}.\text{temp\_match}$
 35:Compute final matching accuracy
@@
 IV Experiments
+IV-A Implementations
 We evaluate the performance of this work from two aspects.
@@
 The loss function follows the same setting in [54].
-Performance metrics include Rank-1, 5, 10, 20, mean Average Precision (mAP), and the Cumulative Matching Characteristic (CMC) curve.
-IV-A Bus ReID Dataset
+Performance metrics include R-1, 5, 10, 20 and mean Average Precision (mAP)
+IV-B Bus ReID Dataset
 To address the current shortage of datasets for occluded re-identification, particularly within public transportation systems, we introduce a new dataset constructed from real-time footage captured over several hours from surveillance cameras mounted at the front and rear doors of buses.
@@
 The dataset also includes images from multiple viewpoints and varying lighting conditions, posing a diverse set of challenges for model training and improving the model’s generalization capabilities in complex environments.
-IV-B Evaluation on ReID Performance
-The performance of TransitReID on traditional ReID tasks is presented in Table I, where it is compared with several state-of-the-art methods, including PGFA [18], HOReID [52], GRL [48], TransReID [51], PFD [34], and CLIP-ReID [53].
-TransitReID achieves the highest accuracy among all comparative methods, improving Rank-1 accuracy by 4.7% and mAP by 8.9%.
-It is noteworthy that PGFA, HOReID, TransReID, and PFD are specifically designed for occluded ReID, which explains their superior performance compared to the vision-language model CLIP-ReID.
-Although our baseline model GRL is not specifically designed for occluded cases, it still outperforms some occluded methods due to its video-based ReID approach that learns temporal correlations.
-Our method not only incorporates temporal continuity information but also selectively focuses on less occluded regions from clear viewpoints to extract useful information.
+IV-C Evaluation on ReID Performance
+The performance of TransitReID on traditional ReID tasks is presented in Table I, where it is compared with several state-of-the-art methods, including PGFA [17], HOReID [57], GRL [54], TransReID [56], PFD [40], CLIP-ReID [58], and TF-CLIP [59].
+TransitReID achieves the highest accuracy among all comparative methods, improving R-1 accuracy by 1.6% and mAP by 1.8%.
+It is noteworthy that PGFA, HOReID, and PFD are specifically designed for occluded ReID, yet our method still achieves better results.
+Although our baseline model GRL is not specifically designed for occluded cases, it still outperforms some occluded methods due to its video-based approach that learns temporal correlations.
+Our method not only incorporates temporal information but also selectively focuses on less occluded regions from clear viewpoints to extract fine-grained information.
 The experimental results show the effectiveness of our method in handling multi-view and complex occlusion scenarios specific to passenger ReID in transit environments.
 TABLE I: Comparison with related methods.
-Matching Rate (%)
-Rank 1 5 10 20 mAP
-TransitReID (Ours) 88.3 96.7 100 100 92.5
-PGFA[18] 60.9 70.2 75.3 80.4 44.3
-HORelD[52] 69.0 75.6 78.2 82.0 58.8
-GRL[48] 75.0 96.7 100 100 83.6
-TransRelD[51] 83.6 87.1 89.1 91.4 78.8
-PFD[34] 78.8 84.0 86.6 89.3 74.6
-CLIP-ReID[53] 55.4 68.4 75.0 82.8 45.7
-Figure 11: Cumulative Matching Characteristic curve of different ReID methods.
-Fig.
-11 presents the CMC curves of TransitReID and its comparative methods from Rank-1 to Rank-20.
-Our method achieves nearly 100% accuracy after Rank-9.
-Notably, the baseline model GRL attains almost identical accuracy with ours after Rank-5, demonstrating its high performance in video-based ReID tasks.
-However, TransitReID significantly outperforms GRL from Rank-1 to Rank-4, validating its effectiveness in suppressing occlusion interference.
+Methods References R-1 R-5 R-10 R-20 mAP
+PGFA[17] ICCV’2019 60.9 70.2 75.3 80.4 44.3
+HORelD[57] CVPR’2020 69.0 75.6 78.2 82.0 58.8
+GRL[54] CVPR’2021 75.0 96.7 100 100 83.6
+TransRelD[56] ICCV’2021 83.6 87.1 89.1 91.4 78.8
+PFD[40] AAAI’2022 78.8 84.0 86.6 89.3 74.6
+CLIP-ReID[58] AAAI’2023 55.4 68.4 75.0 82.8 45.7
+TF-CLIP[59] AAAI’2024 86.7 96.7 98.3 100 90.7
+TransitReID Ours 88.3 96.7 100 100 92.5
+Notably, GRL attains almost identical accuracy with ours after R-5, demonstrating its strong performance in video-based ReID tasks.
+However, TransitReID significantly outperforms GRL in R-1 and mAP, validating its effectiveness in suppressing occlusion interference.
 From another perspective, it enhances the baseline model’s matching capability on ‘hard cases’, where the target gallery person often appears within the top 5 matches but fails to achieve the best match due to difficulties in extracting key features and confusion from very similar samples.
@@
 The score is mapped to colors from dark blue to red, where a higher score indicates higher quality.
-TABLE II: Performance across different numbers of bus stops without / with HSDM.
-Stop ID 5 6 7 8 9 10 11 12 13 14 15
-Rank-1 0.87 / 0.90 0.87 / 0.90 0.87 / 0.91 0.88 / 0.91 0.86 / 0.90 0.88 / 0.92 0.89 / 0.92 0.88 / 0.91 0.90 / 0.92 0.89 / 0.90 0.89 / 0.90
-Rank-5 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
-Rank-10 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
-IV-C Evaluation on Dynamic Transit OD Data Collection
+IV-D Evaluation on Dynamic Transit OD Data Collection
 We randomly assign the 60 passengers in the validation set to several bus stops and randomly determine their boarding and alighting behaviors.
@@
 As a result, the number of gallery samples at each stop in the simulation is overestimated, increasing the matching difficulty but providing a more rigorous evaluation of our method.
+Figure 12: Boarding and alighting simulation with 60 passengers in 10 bus stops.
 Fig.
@@
 Nevertheless, we argue that an accuracy exceeding 90% in the simulation experiment is sufficient to validate the effectiveness of our approach in transit data collection applications.
-Figure 13: Boarding and alighting simulation with 60 passengers in 10 bus stops.
-IV-D Inference Time Analysis on Edge Devices
+IV-E Inference Time Analysis on Edge Devices
 We conduct experiments on an NVIDIA Jetson AGX Orin to simulate the inference time of an edge computing device in a realistic transit scenario.
-The results are presented in Table III.
+The results are presented in Table II.
 Board/Alight indicates the number of passengers boarding or alighting at each stop.
@@
 The last row shows the average inference time per passenger.
-The processing time per passenger is approximately 3 seconds, mainly constrained by the computational performance of the edge device.
+The processing time per passenger is about 3 seconds, mainly constrained by the computational performance of the edge device.
 Nevertheless, We implement multi-threaded programming to enable real-time operation in practical smart transit applications.
 As shown in Fig.
-14, the video thread, connected to the transit surveillance system, captures video frames from the front and rear doors and feeds them into two separate detection threads.
+13, the video thread, connected to the transit surveillance system, captures video frames from the front and rear doors and feeds them into two separate detection threads.
 In these detection threads, real-time processing is achieved due to the high efficiency of the YOLO model.
@@
 The real-time performance of the detection threads ensures smooth operation within the transit system.
-Meanwhile, the featurize thread does not require strict real-time execution—as long as the processing is completed before the next stop.
+Meanwhile, the featurize thread does not require strict real-time execution as long as the processing is completed before the next stop.
 For example, if 10 passengers board the bus at stop $i$, the entire processing should be completed within 30 seconds, before the bus arrives at stop $i+1$.
 Figure 13: Multi-threaded programming.
-TABLE III: Statistics of each stop’s information and processing time.
+TABLE II: Statistics of each stop’s information and processing time.
 Stop ID 1 2 3 4 5
@@
 $T_{F}$ 33.4 30.3 13.2 42.8 36.0
-$T_{M}$ 1.55e-3 1.38e-3 6.30e-3 2.81e-3
-$T_{mean/P}$ 3.37 3.30 3.31 3.00
+$T_{M}$ - 1.55e-3 1.38e-3 6.30e-3 2.81e-3
+$T_{mean/P}$ - 3.37 3.30 3.31 3.00
 Stop ID 6 7 8 9 10
@@
 $T_{mean/P}$ 3.27 3.30 3.06 2.81 3.25
-IV-E Ablation Study
-The effectiveness of the proposed HSDM mechanism is evaluated through simulations with the number of bus stops $S$ ranging from 5 to 15.
+TABLE III: Performance across different numbers of bus stops without / with HSDM.
+Stop ID 5 6 7 8 9 10 11 12 13 14 15
+R-1 0.87 / 0.90 0.87 / 0.90 0.87 / 0.91 0.88 / 0.91 0.86 / 0.90 0.88 / 0.92 0.89 / 0.92 0.88 / 0.91 0.90 / 0.92 0.89 / 0.90 0.89 / 0.90
+R-5 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
+R-10 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
+IV-F Ablation Study
+HSDM
+The effectiveness of the proposed HSDM mechanism is evaluated through simulations with the number of bus stops ranging from 5 to 15.
 Each configuration involves 20 repeated experiments, and the average matching rate across all stops is calculated.
-The results are shown in Table II.
-Without HSDM, the Rank-1 matching rate increases very slightly as the number of bus stops increases.
+The results are shown in Table III.
+Without HSDM, the R-1 matching rate increases very slightly as the number of bus stops increases.
 This is attributed to the reduction in gallery samples at each stop, which enhances the matching success rate.
-With HSDM, the Rank-1 rate increases across all settings, demonstrating the mechanism’s capability to handle low-confidence matches by providing these samples with additional opportunities to achieve the global optimum.
-Notably, with HSDM, the Rank-1 rate remains stable regardless of the number of bus stops.
+With HSDM, the R-1 rate increases across all settings, demonstrating the mechanism’s capability to handle low-confidence matches by providing these samples with additional opportunities to achieve the global optimum.
+Notably, with HSDM, the R-1 rate remains stable regardless of the number of bus stops.
 This stability is achieved through the cold storage mechanism, which ensures that gallery samples at each stop are influenced not only by the number of passengers on board but also by their matching confidence.
-Consequently, the number of matching samples are larger than the actual number of passengers onboard, further validating the robustness of the proposed method under varying numbers of bus stops.
-For Rank-5 and Rank-10, no significant changes are observed with or without HSDM, indicating the high accuracy of our ReID algorithm for simple samples.
+Consequently, the gallery size is larger than the actual number of passengers onboard, further validating the robustness of the proposed method under varying numbers of bus stops.
+For R-5 and R-10, no significant changes are observed with or without HSDM, indicating the high accuracy of our ReID algorithm for simple samples.
+Contributions from Different Components
+We evaluate the contribution of different components on the proposed bus ReID dataset, as shown in Table IV.
+First, we remove one of the three body parts individually, and observe that both mAP and R-1 decrease compared to using the full body.
+However, when the Legs part is missing, R-5 improves, which may be because some hard cases involving legs tend to cause confusion.
+“G” denotes the grading model, which brings improvements of 5.2% in mAP and 10% in R-1.
+SQFA slightly improves R-1 by 0.1% and R-5 by 1.7%.
+TABLE IV: Ablation study on the effectiveness of each component on Bus ReID dataset.
+Head Legs Torso G SQFA mAP R-1 R-5
+84.8 75.0 96.7
+88.1 80.0 98.3
+90.0 85.0 95.0
+89.9 83.3 96.7
+87.2 78.3 95.0
+92.4 88.3 95.0
+92.5 88.3 96.7
 V Conclusion
@@
 As all features are processed and stored locally, passenger privacy is effectively protected.
-Comprehensive evaluations demonstrated the framework’s effectiveness, achieving state-of-the-art performance with 88.3% Rank-1 accuracy on our proposed transit ReID dataset and maintaining on average 90% OD estimation accuracy in real-world transit simulations.
+Comprehensive evaluations demonstrated the framework’s effectiveness, achieving state-of-the-art performance with 88.3% R-1 accuracy on our proposed transit ReID dataset and maintaining on average 90% OD estimation accuracy in real-world transit simulations.
 While real video data are now available, the configuration of bus stops remains a manual process.
-Future work will involve deploying this system in operational public transport vehicles to conduct further tests and refinements, ensuring real-time performance.
+Future work will involve deploying this system in public transport vehicles to conduct further tests and refinements, ensuring real-time performance.
 Moreover, there is a pressing need to develop lighter algorithms based on this research to facilitate real-time execution on edge devices.